Find the midpoint between the two points: (2,2), (6,4)
this is as far as i get (6 - 2)^2 + (4 - 2)^2
4x4 + 2x2
16 + 4 = sqrt20
then i dont know what else to do. can somwone help finish the problem and tell me if i am doing it right

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

Find the midpoint between the two points: (2,2), (6,4)
this is as far as i get (6 - 2)^2 + (4 - 2)^2
4x4 + 2x2
16 + 4 = sqrt20
then i dont know what else to do. can somwone help finish the problem and tell me if i am doing it right
----------------
Sqrt(20) is the distance between the 2 points.
---
To find the midpoint, find the average of x & y separately.
x: (2+6)/2 = 4
y: (2+4)/2 = 3
--> (4,3)

Answer by solver91311(24713)



You are trying to mis-use the distance formula. The measure of the distance between two points is interesting -- but not very -- when you want to find the coordinates of the midpoint. Use the midpoint formulas. The coordinate of the midpoint is the average of the -coordinates of the endpoints. Likewise for the coordinates. In short:

and



where and are the coordinates of the segment end points.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism